OpenStudy (anonymous):

another expression for sin(tan^-1x)

OpenStudy (anonymous):

let : tan y = x/1 sin y = x/ sqrt (x^2 + 1) sin ( sin^-1 y) = y sin ( sin^ -1 x/ sqrt (x^2 + 1)) = x/ sqrt (x^2 + 1)

OpenStudy (anonymous):

where is tan y= x/1 from??

OpenStudy (anonymous):

tan x = 1 x = tan^-1 x

OpenStudy (anonymous):

sorry x = tan^-1 1

OpenStudy (anonymous):

can you please explain it in easier way? i do not understand how it works..

OpenStudy (anonymous):

Watch this video, and if you still need help, we'll jump in.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

i look for the sin(x) because the given equation is sin(arctanx)?

OpenStudy (anonymous):

From diagram, you need to find sin (theta). The guy above seem to be working it algebraically, but you can work it geometrically, just copy it from diagram